OpenStudy (rubylovesjaune):

What is the equation of a line passing through (0, 6) and (3, 0)?

OpenStudy (rubylovesjaune):

Answer choices are Y = -1/2x +6 Y = -1/2x +3 Y = -2x +6 Y = -2x +3

OpenStudy (ijlal):

For finding out and equation of a line we need 2 things 1) A coordinate that passes through the line , here we have 2 so lets use 1 2) Slope or what we call is m

OpenStudy (ijlal):

so we are given 2 points lets find out the slope first